More than one option could be correct.Question 3: (4 pts) Fill in the circle next to each vector-valued function that parameterizes the ellipse 4.x2 + y2 = 4 in th
Answer #1

3 (t) = 444) Ĉ + y(t) {(t) = < Kitt), y t)) Ellipse 4x+y+= 4 . LHS = 4x++y2 R15 = 4 LMS means Left hand side RHS means RightX(t), yt) of {(t)= < htt), yt) makes LHS = RHS ® {(t) = <t, J4-47²) 1 + = + . g(t) = J4-4+2 . Putting these in LHS LHS= 48²42Lus= 4*+ 4-44 LHS= 4 LHS=KHS so; option A is correct (B) 32(t)=(2cost, seint> 744)= 2lost y(t) = sint peeting these in LasLHS= 4x²+ y2 215 = 4(2cost)? + Esentj2 Ius = 414 cost) + sin til LH S=16 costt sinit Lus=15 cost + cos²tt sint LHS=15cost(senza . tcostA=1] so, LKS & RHS. so, option B is corong e az(t) = 5 cost, 2 sint) ut) = cost y(t) = sint putting these in LH2 LHS = 4(cost)+(2 sin t) LUS= 4 cost + 4 sint LHS = 4 (cost tsinnt) LKS= 4 tsin At Costa -17 LHS= RHS. so, option c is corЧЕ) = 2 (26) рина Илих лh Lux LHS = 4x+y2 Lus = 4 (ct) + 2 даn(21)] LHS= 4 cos²t + 4 sin 2tLHS = 4 cost t 4 sin2t LHs= 4 [cost + sen 22t] Lms=4[cost + [sen 2t)) L95 = 4 [ cost + (Aseint cost)) LHs = 4 [ coo*t + 4 seso, option (A) and option() are correct. VA) Rit)= <t,d4-46) ver) R35+) =< cost, 2sist>